OpenStudy (sherribelle):

2P=Pe^rt. a. Solve the equation for t in terms of r. b. How long does it take for an investment to double in value compounding continuously if the annual interest rate is 7%? Round your answer to the nearest tenth of a year. c. Solve the equation for r in terms of t. d. What interest rate is necessary for an investment to double in value in 7 years if interest is compounded continuously? Round your answer to the nearest tenth of a percent.

OpenStudy (sherribelle):

I got some of it, however, I don't know if it's correct.

OpenStudy (sherribelle):

@matthew120

OpenStudy (matthew120):

hmmm

OpenStudy (matthew120):

what one do u think it is

OpenStudy (sherribelle):

I don't know if I'm doing it correctly, and it's 4 different steps, not multiple choice.

OpenStudy (matthew120):

i bleive its a not 100% ok

OpenStudy (matthew120):

believe

OpenStudy (sherribelle):

so far what i got is: a) t = ln(2)/r + ln(P)/r b) 14.3 c) r = ln(2)/t + ln(P)/t d) 14%

OpenStudy (sherribelle):

I don't know if any of that is correct.

jhonyy9 (jhonyy9):

rt - every two are exponents ?

jhonyy9 (jhonyy9):

or just r is exponent ?

OpenStudy (sherribelle):

rt are exponents

jhonyy9 (jhonyy9):

2P = P*e^(rt) in this way ?

OpenStudy (sherribelle):

yes and e is actually a base of natural logarithms. it equals like 2.71828. I have to look up the exact answer

jhonyy9 (jhonyy9):

ok Sherri but like a first step you can divide both sides by P - yes ?

OpenStudy (sherribelle):

you could

jhonyy9 (jhonyy9):

bc there is P on both sides and you need to solve it for t in term of r ok. ?

OpenStudy (sherribelle):

Okay, so it cancels the P?

OpenStudy (sherribelle):

Wait one P has 2

jhonyy9 (jhonyy9):

after this 2 = e^(rt) make ln on both sides 2lne = lne^(rt) bc. lne = 1 and so 2lne = 2 lne^2 = lne^(rt) than you see lne are common so this will be equal than these exponents will be equal - so 2 = rt t = 2/r

OpenStudy (sherribelle):

so then would you replace r with 0.07 in t = 2/r?

jhonyy9 (jhonyy9):

or r = 2/t

jhonyy9 (jhonyy9):

i think yes

jhonyy9 (jhonyy9):

for this 7%

OpenStudy (sherribelle):

so t would be 28.6 and r would be 28%?

jhonyy9 (jhonyy9):

i think it so yes

OpenStudy (sherribelle):

hmmm okay

jhonyy9 (jhonyy9):

good luck bye bye